# PoisonedFL: Model Poisoning Attacks to Federated Learning via Multi-Round Consistency

### v1

@@ 5. PoisonedFL
Figure 2. Illustration of the global-model evolution in three training rounds under existing attacks and PoisonedFL. The attack effect self-cancels in existing attacks, while PoisonedFL consistently moves the global model along the same direction.
5.1. Overview
@@
5.2. Formulating an Optimization Problem
@@
(2) $\begin{gathered}\max_{\bm{g}_{i}^{t},i\in[n+1,n+m],t\in[1,T]}\left\|\sum_{t=1}^{T}\bm{g}^{t}\right\|\\ s.t.\ \text{sign}(\sum_{t=1}^{T}\bm{g}^{t})=\bm{s},\end{gathered}$
@@
(3) $\max_{\bm{g}_{i}^{t},i\in[n+1,n+m]}\left\|\bm{g}^{t}\right\|,\ s.t.\ \text{sign}(\bm{g}^{t})=\bm{s}.$
Since the aggregated model updates across training rounds have the same sign vector $\bm{s}$, their magnitudes add up in the total aggregated model update, leading to a bad global model.
5.3. Solving the Optimization Problem
@@
(4) $\bm{g}_{i}^{t}=\bm{k}^{t}\odot\bm{s},\forall i\in[n+1,n+m],\forall t\in[1,T],$
@@
Recall that the objective function of our per-round optimization problem is to maximize the magnitude of the aggregated model update $\left\|\bm{g}^{t}\right\|$. One naive way to achieve this objective is to select a large magnitude vector $\bm{k}^{t}$. However, such a magnitude vector $\bm{k}^{t}$ may be filtered out by the server when a defense (e.g., Multi-Krum (Blanchard et al., 2017), FLTrust (Cao et al., 2021a), or FLAME (Nguyen et al., 2022)) is deployed, leading to suboptimal attack effectiveness as our experiments in Section 6 show. Therefore, our goal is to maximize the magnitude vector $\bm{k}^{t}$ while avoiding being filtered out by an unknown defense that may be deployed by the server. To achieve this goal, our key idea is to pick $\bm{k}^{t}$ that has a large magnitude but is also similar to the model updates on the genuine clients. In particular, we explicitly decompose the magnitude vector $\bm{k}^{t}$ into the following form:
(5) $\bm{k}^{t}=\lambda^{t}\cdot\bm{v}^{t},$
@@
(6) $\bm{v}^{t}=\frac{\left|\bm{g}^{t-1}-\frac{\left\|\bm{g}^{t-1}\right\|}{\left\|\bm{k}^{t-1}\odot\bm{s}\right\|}\cdot\bm{k}^{t-1}\odot\bm{s}\right|}{\left\|\bm{g}^{t-1}-\frac{\left\|\bm{g}^{t-1}\right\|}{\left\|\bm{k}^{t-1}\odot\bm{s}\right\|}\cdot\bm{k}^{t-1}\odot\bm{s}\right\|},$
@@
(7) $\lambda^{t}=c^{t}\cdot\left\|\bm{w}^{t-1}-\bm{w}^{t-2}\right\|,$
@@
6. Evaluation
6.1. Experimental Setup
6.1.1. Datasets
We use five datasets from different domains, including MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), Purchase (pur, 2021), CIFAR-10 (Krizhevsky et al., 2009), and FEMNIST (Caldas et al., 2018). Details are deferred to Appendix B.1.
6.1.2. Evaluation Metric
@@
6.1.3. Defenses and Compared Attacks
We evaluate eight state-of-the-art defenses, including six Byzantine-robust aggregation rules (Multi-Krum (Blanchard et al., 2017), Median (Yin et al., 2018), Trimmed Mean (Yin et al., 2018), Norm Bound (Sun et al., 2019), FLTrust (Cao et al., 2021a), and FLAME (Nguyen et al., 2022)), one provably robust defense (FLCert (Cao et al., 2022)) and one malicious clients detection method (FLDetector (Zhang et al., 2022)). We also consider the non-robust FedAvg (McMahan et al., 2017) as a baseline. We compare our PoisonedFL with seven attacks, including five attacks (Fang (Fang et al., 2020), Opt. Fang (Shejwalkar and Houmansadr, 2021), LIE (Baruch et al., 2019), Min-Max (Shejwalkar and Houmansadr, 2021), and Min-Sum (Shejwalkar and Houmansadr, 2021)) that require genuine clients’ information and two attacks (Random (Cao and Gong, 2022) and MPAF (Cao and Gong, 2022)) that do not require such information. Details are deferred to Appendix B.3 and B.4.
6.1.4. Evaluation Metric
6.1.5. FL Settings
By default, we assume $1,200$ genuine clients and 20% of fake clients, i.e., the number of fake clients is 20% of the number of genuine clients. In each training round, the server selects a random subset of 10% of clients to participate in training. A genuine client trains its local model for 1 local epoch in a training round using SGD. Due to limited space, more FL settings for the five datasets are detailed in Appendix B.2. We also study the impact of different FL settings (fraction of fake clients, non-IID degree, local training epochs, and participation rate) in Section 6.3. Our PoisonedFL has parameters $c^{t}$, $e$, and $\beta$. We set $c^{0}=8$, $e=50$, and $\beta=0.7$, and we do not decrease $c^{t}$ to be smaller than $0.5$ since a too small $c^{t}$ is ineffective. We conduct our experiments using 16 GeForce RTX 3090 GPUs.
Table 1. Testing error rate (%) of the global model under different attacks and defenses for different datasets. We assume Fang, Opt. Fang, LIE, Min-Max, and Min-Sum have access to the model updates on all genuine clients, and Fang and Opt. Fang further have access to the aggregation rule, which give advantages to these attacks. Random, MPAF, and PoisonedFL do not have access to these knowledge. In each row, we highlight in bold the highest testing error rate and those no better than random guessing.
@@ 6.2. Main Results
(a) Multi-Krum
(b) Median
(c) TrMean
(d) Norm Bound
(e) FLTrust
(f) FLAME
(g) FLCert
(h) FLDetector
Figure 3. Testing error rate of the global model as a function of the fraction of fake clients (%) under different defenses and attacks.
(a) Multi-Krum
(b) Median
(c) TrMean
(d) Norm Bound
(e) FLTrust
(f) FLAME
(g) FLCert
(h) FLDetector
Figure 4. Testing error rate of the global model as a function of the degree of non-IID under different defenses and attacks.
@@ 6.3. Impact of FL Settings
(h) FLDetector
Figure 5. Testing error rate of the global model as a function of local training epochs under different defenses and attacks.
(a) Multi-Krum
@@
(h) FLDetector
Figure 6. Testing error rate of the global model as a function of the participation rate under different defenses and attacks.
@@
Impact of the degree of non-IID:  In our experiments, we use a parameter $q$ to control the degree of non-IID among the genuine clients’ local training data, where $0.1\leq q\leq 1$ (Fang et al., 2020). A larger $q$ indicates a larger degree of non-IID, where $q=0.1$ means the genuine clients’ local training data is IID and $q=1$ indicates highly non-IID (i.e., each class of data is distributed among a subset of clients). Fig. 4 shows the testing error rate of the global model as $q$ increases from 0.1 to 1 under different attacks and defenses. We observe that PoisonedFL consistently achieves high testing error rates (i.e., random guessing) across different settings of $q$ under different defenses. However, it is harder for existing attacks to compromise the defenses when $q$ is smaller, e.g., $q=0.1$. This is because, in such cases, the genuine model updates are less diverse, leaving less room for existing attacks to craft malicious model updates. We note that existing attacks are substantially less effective than PoisonedFL under Multi-Krum and FLTrust even if $q=1.0$.
@@
6.4. Impact of Different PoisonedFL Variants
@@
Figure 7. (a) Comparing the two variants of setting the unit magnitude vector in PoisonedFL. (b) Comparing the two variants of setting the scaling factor in PoisonedFL.
@@
  • Adaptive scaling factor: This variant, adopted by PoisonedFL, leverages how the global model changes in the previous training round to set the scaling factor (i.e., Equation 7) to avoid that the malicious model updates are filtered out by the server.
@@
7. Countermeasures and Adaptive Attacks
Refer to captionRefer to caption
(a) Median
Refer to captionRefer to caption
(b) TrMean
Refer to captionRefer to caption
(c) Norm Bound
Figure 8. Detection accuracy of GMM-Sign and testing error rate of the global model under PoisonedFL-Adapt attack as the flipping rate $\alpha$ varies when the server uses different aggregation rules.
Refer to captionRefer to caption
(a) Median
Refer to captionRefer to caption
(b) TrMean
Refer to captionRefer to caption
(c) Norm Bound
Figure 9. Detection accuracy of GMM-Magnitude and testing error rate of the global model under PoisonedFL-Adapt attack as the noise scaling factor $\gamma$ varies when the server uses different aggregation rules.
@@ 7. Countermeasures and Adaptive Attacks
In our tailored detection based defenses, the server extracts a feature for each client in each training round and then divides the clients into two clusters via fitting a two-component Gaussian Mixture Model (GMM) (Permuter et al., 2006) for the features. We detect a particular cluster (e.g., with the smaller mean value) as fake clients. After detecting fake clients, the server removes them and re-trains a global model using the remaining clients. Note that if the distance between the mean values of the two clusters is smaller than the intra-cluster standard deviations of both clusters, no clients are detected as fake and the training continues since the difference between the two clusters may simply be caused by randomness.
@@ 7.1. GMM-Sign
(8) $x_{i}^{t}=\sum_{t^{\prime}=t-N+1}^{t}\sum_{j=1}^{d}\mathbb{I}(\text{sign}(\bm{g}_{i}^{t^{\prime}}[j])\neq\text{sign}(\bm{g}_{i}^{t^{\prime}-1}[j])),$
@@
(9) $\begin{split}\bm{s}^{t}_{i}[j]&=\begin{cases}-\bm{s}{[j]}&\text{with probability }\alpha\\ \bm{s}{[j]}&\text{with probability }1-\alpha\end{cases},\\ \bm{k}^{t}_{i}[j]&=\begin{cases}\epsilon&\text{if }\bm{s}^{t}_{i}[j]=-\bm{s}{[j]}\\ \bm{k}^{t}{[j]}&\text{otherwise }\end{cases},\end{split}$
@@ 7.2. GMM-Magnitude
(10) $x_{i}^{t}=\sum_{t^{\prime}=t-N+1}^{t}{{\sum_{j\in\mathbb{N}_{i}^{t^{\prime}}}\frac{\left\|\bm{g}_{i}^{t^{\prime}}-\bm{g}_{j}^{t^{\prime}}\right\|^{2}}{m-1}}},$
@@
(11) $\bm{k}^{t}_{i}=\bm{k}^{t}+\gamma\cdot\frac{||\bm{k}^{t}||}{||\bm{\epsilon}_{i}||}\cdot\bm{\epsilon}_{i},$
@@
7.3. Normalization-based Defense
@@ 8. Discussion and Limitations
Reconstructing training data from a model is known as model inversion (Fredrikson et al., 2015). Previous work (Pi et al., 2023) shows that the global-model trajectory in multiple training rounds (especially early ones) of FL can be used to reconstruct synthetic data that is similar to the clients’ genuine local training data. We follow the setting and generation strategy in (Pi et al., 2023), which is the state-of-the-art reconstruction method in FL, to generate synthetic data for MNIST and FashionMNIST. We use MNIST and FashionMNIST because they are simpler datasets and easier to reconstruct, giving advantages to these attacks. Specifically, we assume no attacks in the first $40$ rounds and use the corresponding clean global models to reconstruct synthetic data. Table 2 shows the testing error rates of the classifiers trained on the synthetic data and evaluated on the genuine testing datasets of MNIST and FashionMNIST. The results show that the classifiers have decent testing error rates and thus the synthetic data mimic the genuine data well, which is consistent with prior work (Pi et al., 2023).
Table 2. Testing error rate of a classifier trained on synthetic data and evaluated on genuine testing data.
@@
Table 3. Testing error rates of attacks using synthetic data on fake clients (i.e., “Syn. Data + ”) and local models of all genuine clients (i.e., “Genuine + ”), as well as our PoisonedFL that does not require synthetic data nor genuine local models.
@@
Cross-silo FL:  PoisonedFL primarily targets cross-device FL (Kairouz et al., 2021), in which the clients are end-user devices such as smartphones. An attacker can inject fake clients into a cross-device FL system due to the open nature of the system. In particular, any client including fake ones can participate in the FL system. We acknowledge that it is harder to apply both PoisonedFL and existing attacks to cross-silo FL. Specifically, the clients are often a small number of verified institutions such as banks and hospitals in cross-silo FL. These institutions often go through certain verification processes before jointly training a model using FL. As a result, it is harder for an attacker to compromise genuine clients/institutions or inject fake clients/institutions in such a cross-silo FL system. We believe it is an interesting future work to identify a realistic threat model for cross-silo FL that considers the unique challenges and constraints.
9. Conclusion

### v2

@@ 5.1 Overview
Figure 2: Illustration of the global-model evolution in three training rounds under existing attacks and PoisonedFL. The attack effect self-cancels in existing attacks, while PoisonedFL consistently moves the global model along the same direction.
@@
5.2 Formulating an Optimization Problem
@@
$\begin{gathered}\max_{\bm{g}_{i}^{t},i\in[n+1,n+m],t\in[1,T]}\left\|\sum_{t=1}^{T}\bm{g}^{t}\right\|\\ s.t.\ \text{sign}(\sum_{t=1}^{T}\bm{g}^{t})=\bm{s},\end{gathered}$ (2)
@@
$\max_{\bm{g}_{i}^{t},i\in[n+1,n+m]}\left\|\bm{g}^{t}\right\|,\ s.t.\ \text{sign}(\bm{g}^{t})=\bm{s}.$ (3)
Since the aggregated model updates across rounds have the same sign vector $\bm{s}$, their magnitudes add up in the total aggregated model update, leading to a bad global model.
5.3 Solving the Optimization Problem
@@
$\bm{g}_{i}^{t}=\bm{k}^{t}\odot\bm{s},\forall i\in[n+1,n+m],\forall t\in[1,T],$ (4)
@@
Recall that the objective function of our per-round optimization problem is to maximize the magnitude of the aggregated model update $\left\|\bm{g}^{t}\right\|$. One naive way to achieve this objective is to select a large magnitude vector $\bm{k}^{t}$. However, such a magnitude vector $\bm{k}^{t}$ may be filtered out by the server when a defense (e.g., Multi-Krum [12], FLTrust [13], or FLAME [15]) is deployed, leading to suboptimal attack effectiveness as our experiments in Section 6 show. Therefore, our goal is to maximize the magnitude vector $\bm{k}^{t}$ while avoiding being filtered out by an unknown defense that may be deployed by the server. To achieve this goal, our key idea is to pick $\bm{k}^{t}$ that has a large magnitude but is also similar to the model updates on the genuine clients. In particular, we explicitly decompose the magnitude vector $\bm{k}^{t}$ into the following form:
$\bm{k}^{t}=\lambda^{t}\cdot\bm{v}^{t},$ (5)
@@
$\bm{v}^{t}=\frac{\left|\bm{g}^{t-1}-\frac{\left\|\bm{g}^{t-1}\right\|}{\left\|\bm{k}^{t-1}\odot\bm{s}\right\|}\cdot\bm{k}^{t-1}\odot\bm{s}\right|}{\left\|\bm{g}^{t-1}-\frac{\left\|\bm{g}^{t-1}\right\|}{\left\|\bm{k}^{t-1}\odot\bm{s}\right\|}\cdot\bm{k}^{t-1}\odot\bm{s}\right\|},$ (6)
@@
$\lambda^{t}=c^{t}\cdot\left\|\bm{w}^{t-1}-\bm{w}^{t-2}\right\|,$ (7)
@@
6 Evaluation
6.1 Experimental Setup
6.1.1 Datasets
We use the following five datasets from different domains.
MNIST [33]:  MNIST is a 10-class handwritten digits classification dataset. It comprises 60,000 training examples and 10,000 testing examples. In FL, the training data is typically not independently and identically distributed (non-IID) among clients. Following [9, 10], we distribute the training examples among clients based on a non-IID degree of $q=0.5$ by default. We train a convolutional neural network (CNN) for MNIST dataset. The architecture of CNN is shown in Table 9 in Appendix.
FashionMNIST [34]:  FashionMNIST is a benchmark dataset of Zalando’s article images, which contains 60,000 training examples and 10,000 testing examples. For FashionMNIST, the degree of non-IID is also set to $q=0.5$ by default. We use the same CNN architecture for FashionMNIST dataset as employed in MNIST dataset.
Purchase [35]:  Purchase is a 100-class customer purchase style prediction dataset. Each input consists of 600 binary features. Following [9], we split the total 197,324 examples into 180,000 training examples and 17,324 testing examples. Following [9], we evenly distribute the training examples to clients and train a fully connected neural network.
CIFAR-10 [36]:  CIFAR-10 is 10-class color image classification dataset. It comprises 50,000 training examples and 10,000 testing examples. To consider a different degree of non-IID, we set $q=0.2$ for CIFAR-10 dataset. We train a CNN for CIFAR-10, whose architecture is shown in Table 10.
FEMNIST [37]:  FEMNIST is a 62-class classification dataset. The dataset is already distributed among 3,550 clients with a total of 805,263 examples. We randomly sample 1,200 clients. We train a CNN for FEMNIST, whose architecture is shown in Table 11.
6.1.2 Defenses and Compared Attacks
We evaluate eight state-of-the-art defenses, including six Byzantine-robust aggregation rules, i.e., Multi-Krum (M-Krum) [12], Median [17], Trimmed Mean (TrMean) [17], Norm Bound (N-Bound) [16], FLTrust [13], and FLAME [15], one provably robust defense, i.e., FLCert [14], and one malicious clients detection method, i.e., FLDetector (FLDet) [18]. We also consider the non-robust FedAvg [2] as a baseline. We compare our PoisonedFL with seven attacks, including five attacks (Fang [10], Opt. Fang [11], LIE [8], Min-Max [11], and Min-Sum [11]) that require genuine clients’ information and two attacks (Random [9] and MPAF [9]) that do not require such information. Details are deferred to Appendix B.2 and B.3.
6.1.3 Evaluation Metric
@@
6.1.4 FL Settings
By default, we assume $1,200$ genuine clients and 20% of fake clients, i.e., the number of fake clients is 20% of the number of genuine clients. In each training round, the server selects a random subset of 10% of clients to participate in training. A genuine client trains its local model for 1 local epoch in a training round using SGD. Due to limited space, more FL settings for the five datasets are detailed in Appendix B.1. We also study the impact of different FL settings (fraction of fake clients, non-IID degree, local training epochs, and participation rate) in Section 6.3. Our PoisonedFL has parameters $c^{t}$, $e$, and $\beta$. We set $c^{0}=8$, $e=50$, and $\beta=0.7$, and we do not decrease $c^{t}$ to be smaller than $0.5$ since a too small $c^{t}$ is ineffective.
Table 1: Testing error rate (%) of the global model under different attacks and defenses for different datasets. We assume Fang, Opt. Fang, LIE, Min-Max, and Min-Sum have access to the model updates on all genuine clients, and Fang and Opt. Fang further have access to the aggregation rule, which give advantages to these attacks. Random, MPAF, and PoisonedFL do not have access to these knowledge. In each row, we highlight in bold the highest testing error rate and those no better than random guessing.
@@ 6.3 Impact of FL Settings
(h) FLDet
Figure 3: Testing error rate of the global model as a function of the fraction of fake clients (%) under different defenses and attacks.
(a) M-Krum
@@
(h) FLDet
Figure 4: Testing error rate of the global model as a function of the degree of non-IID under different defenses and attacks.
(a) M-Krum
(b) Median
(c) TrMean
(d) N-Bound
(e) FLTrust
(f) FLAME
(g) FLCert
(h) FLDet
Figure 5: Testing error rate of the global model as a function of local training epochs under different defenses and attacks.
(a) M-Krum
(b) Median
(c) TrMean
(d) N-Bound
(e) FLTrust
(f) FLAME
(g) FLCert
(h) FLDet
Figure 6: Testing error rate of the global model as a function of the participation rate under different defenses and attacks.
@@
Impact of the degree of non-IID:  In our experiments, we use a parameter $q$ to control the degree of non-IID among the genuine clients’ local training data, where $0.1\leq q\leq 1$ [10]. A larger $q$ indicates a larger degree of non-IID, where $q=0.1$ means the genuine clients’ local training data is IID and $q=1$ indicates highly non-IID (i.e., each class of data is distributed among a subset of clients). Fig. 4 shows the testing error rate of the global model as $q$ increases from 0.1 to 1 under different attacks and defenses. We observe that PoisonedFL consistently achieves high testing error rates (i.e., random guessing) across different settings of $q$ under different defenses. However, it is harder for existing attacks to compromise the defenses when $q$ is smaller, e.g., $q=0.1$. This is because, in such cases, the genuine model updates are less diverse, leaving less room for existing attacks to craft malicious model updates. We note that existing attacks are substantially less effective than PoisonedFL under Multi-Krum and FLTrust even if $q=1.0$.
@@
Figure 7: (a) Comparing the two variants of setting the unit magnitude vector in PoisonedFL. (b) Comparing the two variants of setting the scaling factor in PoisonedFL.
6.4 Impact of Different PoisonedFL Variants
@@
  • Adaptive scaling factor: This variant, adopted by PoisonedFL, leverages how the global model changes in the previous training round to set the scaling factor (i.e., Equation 7) to avoid the malicious model updates are filtered out by the server.
@@
7 Countermeasures and Adaptive Attacks
@@
In our tailored detection based defenses, the server extracts a feature for each client in each training round and then divides the clients into two clusters via fitting a two-component Gaussian Mixture Model (GMM) [38] for the features. We detect a particular cluster (e.g., with the smaller mean value) as fake clients. After detecting fake clients, the server removes them and re-trains a global model using the remaining clients. Note that if the distance between the mean values of the two clusters is smaller than the intra-cluster standard deviations of both clusters, no clients are detected as fake and the training continues since the difference between the two clusters may simply be caused by randomness.
@@ 7.1 GMM-Sign
$x_{i}^{t}=\sum_{t^{\prime}=t-N+1}^{t}\sum_{j=1}^{d}\mathbb{I}(\text{sign}(\bm{g}_{i}^{t^{\prime}}[j])\neq\text{sign}(\bm{g}_{i}^{t^{\prime}-1}[j])),$ (8)
@@
$\begin{split}\bm{s}^{t}_{i}[j]&=\begin{cases}-\bm{s}{[j]}&\text{with probability }\alpha\\ \bm{s}{[j]}&\text{with probability }1-\alpha\end{cases},\\ \bm{k}^{t}_{i}[j]&=\begin{cases}\epsilon&\text{if }\bm{s}^{t}_{i}[j]=-\bm{s}{[j]}\\ \bm{k}^{t}{[j]}&\text{otherwise }\end{cases},\end{split}$ (9)
@@
(a) Median
(b) TrMean
(c) Norm Bound
Figure 8: Detection accuracy of GMM-Sign and testing error rate of the global model under PoisonedFL-Adapt attack as the flipping rate $\alpha$ varies when the server uses different aggregation rules.
(a) Median
(b) TrMean
(c) Norm Bound
Figure 9: Detection accuracy of GMM-Magnitude and testing error rate of the global model under PoisonedFL-Adapt attack as the noise scaling factor $\gamma$ varies when the server uses different aggregation rules.
@@ 7.2 GMM-Magnitude
$x_{i}^{t}=\sum_{t^{\prime}=t-N+1}^{t}{{\sum_{j\in\mathbb{N}_{i}^{t^{\prime}}}\frac{\left\|\bm{g}_{i}^{t^{\prime}}-\bm{g}_{j}^{t^{\prime}}\right\|^{2}}{m-1}}},$ (10)
@@
$\bm{k}^{t}_{i}=\bm{k}^{t}+\gamma\cdot\frac{||\bm{k}^{t}||}{||\bm{\epsilon}_{i}||}\cdot\bm{\epsilon}_{i},$ (11)
@@
7.3 Normalization-based Defense
@@ 8 Discussion and Limitations
Reconstructing training data from a model is known as model inversion [39]. Previous work [40] shows that the global-model trajectory in multiple training rounds (especially early ones) of FL can be used to reconstruct synthetic data that is similar to the clients’ genuine local training data. We follow the setting and generation strategy in [40], which is the state-of-the-art reconstruction method in FL, to generate synthetic data for MNIST and FashionMNIST. We use MNIST and FashionMNIST because they are simpler datasets and easier to reconstruct, giving advantages to these attacks. Specifically, we assume no attacks in the first $40$ rounds and use the corresponding clean global models to reconstruct synthetic data. Table 2 shows the testing error rates of the classifiers trained on the synthetic data and evaluated on the genuine testing datasets. The results show that the classifiers have decent testing error rates and thus the synthetic data mimic the genuine data well, which is consistent with prior work [40].
Table 2: Testing error rate of a classifier trained on synthetic data and evaluated on genuine testing data.
@@
Table 3: Testing error rates of attacks using synthetic data on fake clients (i.e., “Syn. Data + ”) and local models of all genuine clients (i.e., “Genuine + ”), as well as our PoisonedFL that does not require synthetic data nor genuine local models.
@@
Cross-silo FL:  PoisonedFL primarily targets cross-device FL [4], in which the clients are end-user devices such as smartphones. An attacker can inject fake clients into a cross-device FL system due to the open nature of the system. In particular, any client including fake ones can participate in the FL system. We acknowledge that it is harder to apply both PoisonedFL and existing attacks to cross-silo FL. Specifically, the clients are often a small number of verified institutions such as banks and hospitals in cross-silo FL. These institutions often go through certain verification processes before jointly training a model using FL. As a result, it is harder for an attacker to compromise genuine clients/institutions or inject fake clients/institutions in such a cross-silo FL system. We believe it is an interesting future work to identify a realistic threat model for cross-silo FL that considers the unique challenges and constraints.
9 Conclusion